# Leveraging MLLM Embeddings and Attribute Smoothing for Compositional Zero-Shot Learning

### v1

@@ 4.3 Ablation Study
Impact of word embeddings. Our work leverages the last hidden states of LLaVA v1.5 (${LLaVA}_{lhs}$) as word embeddings, while Word2Vec [21] and Fasttext [3] are the most common word embeddings for MIT-States in previous works. In Table 3, based on three models: SCEN [13], CANet [42] and TRIDENT, we compare the performance of employing the last hidden states of LLaVA v1.5 and the sum of Word2Vec and Fasttext (ft+w2v), respectively. The results indicate that the last hidden states of MLLM capture more complex multimodal semantic information than ordinary word embeddings.
For details on the impact of hyperparameters, including the number of visual features and the label smoothing factor, please refer to Appendix E.
@@ 4.4 Qualitative Analysis
Inspired by [8], we use TRIDENT to conduct both image-to-text retrieval and text-to-image retrieval experiments on the three datasets. We first consider image-to-text retrieval, shown in Figure 3(a). The first two rows display successful cases, while the last row presents failure cases. And the cases shown in these three columns are drawn from the three datasets, respectively. Given an image, such as the image of burnt house, we extract its visual features by image embedder and retrieve the top-5 closest composed pair embeddings of compositions. For successful cases, such as the image labeled burnt house, we notice that the top four predictions can both describe logs burning on fire in the image. In terms of the image labeled green leaf, another successful case, the predicted attributes can also describe leaf, which is thanks to attribute smoothing learning more attributes for an object. For the failure cases, such as the image labeled multicolored teddy-bear, the model focuses on the main orange bear and neglects the multicolored background, which is attributed to FAA modules.
We then consider text-to-image retrieval. Successful cases are shown in Figure 3(b), while failure cases are shown in Figure 3(c). Given a text composition, we embed it and retrieve the top-5 closest images. We observe that the retrieved images of peeled orange are definitely correct. However, the retrieved images of green grapes are all wrong. This is due to the fact that the training images of green grapes in C-GQA dataset are almost filled with a single grape, making it difficult for the model to capture the contour features of a bunch of green grapes. The image-to-text and text-to-image retrieval experiments confirm that our model effectively projects visual features and word embeddings into a unified space.
@@ 5 Conclusion
In this work, we propose a novel framework termed TRIDENT to address the challenging CZSL task. First, we leverage feature adaptive aggregation modules to mitigate the impact of background, and utilize learnable condition masks to capture multi-granularity features for attribute-object disentanglement. In addition, we exploit the last hidden states of MLLM to replace ordinary word embeddings, as they can capture complex multimodal semantic information. Moreover, we leverage LLM to generate auxiliary attributes and perform attribute smoothing to diminish overconfidence of models in seen compositions, which enables models to generalize to unseen compositions better. Extensive experiments have been conducted on three challenging datasets, and the results demonstrate the effectiveness of TRIDENT. In the future, we plan to extend our method to harness the powerful capabilities of LLMs, MLLMs, and CLIP to more effectively address the CZSL task.
Appendix
@@ Appendix A Auxiliary Attributes Generation by LLM and Post-process
As mentioned before, we leverage GPT-3.5 to generate some auxiliary attributes for attribute smoothing. The auxiliary attributes are generated based on the contextual composition of the object and its attribute, such as ripe apple or sliced apple. The model takes into account the unique characteristics that arise from the combination of attribute and the object’s context. This ensures that the auxiliary attributes accurately capture the nuances of the specific composition, rather than general object-level attributes. Since the generation ability of LLM is affected by many factors, in this section, we first explore the impact of different prompts on LLM-generated content. Then we study on the influence of the number of auxiliary attributes $t$. In addition, we introduce post-processing of generated text to eliminate potential noise.
@@
Impact of prompt input to LLM. Users typically interact with LLMs through prompts to request answers to questions, generate text, complete tasks, and more. The model generates text based on the provided prompt, striving to meet the user’s requirements [37]. Therefore, the good design of prompt is significant for stimulating knowledge of LLMs, which enables them to better follow our instructions and generate auxiliary attributes with high quality. We first design some prompts with different style, then input them into GPT-3.5 and observe the quality of generated attributes. Some prompt examples on the composition browned beef and ancient building are shown in Table 4 and Table 5, respectively.
As shown in Table 4, the prompt without ”the visual feature of …” may cause the model to produce adjectives that are not specific but generic, such Savory and Rich. In both Table 4 and Table 5, the prompts starting with ”Now I give you…”, compared to those starting with ”Please give me …”, result in a weaker instruction following ability of the model. Therefore, we choose the prompt: ”Please give me five adjectives that can describe the visual feature of a photo of a/an … well.”
Composition Train Validation Test
Dataset $|\mathcal{A}|$ $|\mathcal{O}|$ $|\mathcal{A}\times\mathcal{O}|$ $|\mathcal{C}_{s}|$ $|\mathcal{X}|$ $|\mathcal{C}_{s}|$ $|\mathcal{C}_{u}|$ $|\mathcal{X}|$ $|\mathcal{C}_{s}|$ $|\mathcal{C}_{u}|$ $|\mathcal{X}|$
MIT-States [11] 115 245 28175 1262 30338 300 300 10420 400 400 12995
C-GQA [25] 413 674 278362 5592 26920 1252 1040 7280 888 923 5098
VAW-CZSL [38] 440 541 238040 11175 72203 2121 2322 9524 2449 2470 10856
Table 6: Summary statistics of the datasets used in our experiments.
Impact of the number of auxiliary attributes $t$. In Table 4, we observe that the generated attributes describe the compositions to varying degrees, with later items in the sequence being less relevant generally. Therefore, we study on the influence of the number of auxiliary attributes $t$.
@@
Table 7 and Table 8 show the generated text using different $t$ of compositions large garden and young girl. The results demonstrate that the greater the number, the more generic adjectives with irrelevant information are included, for example, Captivating is generated for both compositions. In addition, with $t$ increasing, the noise in the generated text due to the uncertainty of the model about the given image grows. The young girl may not be happy, yet the model fails to find ten words to describe her, so it has to guses. Therefore, we set $t$ to 3, this minimizes the general adjectives and noise while retaining useful information.
@@
Post-processing of generated text. GPT-3.5 generates a segment of text, which we need to process into multiple useful words by exploiting regular expressions. However, the auxiliary attributes generated by LLM may contain the attribute of the input composition, for example, generating ancient for ancient building. At this point, we reuse the model to generate $t+1$ adjectives for this composition and select three adjectives that are not the attribute of the input composition.
Appendix B Obtainment of The Last Hidden States of MLLM
@@
We input the attribute (object) word into LLaVA v1.5, which first tokenizes the word into z tokens. These tokens pass through all attention blocks in the MLLM, ultimately generating z embeddings of dimension $d_{m}$ after the last block, named the last hidden states. Subsequently, we apply average pooling to these z embeddings of dimension $d_{m}$ to obtain a $d_{m}$-dimensional embedding that represents the attribute. Since the last hidden states are designed to generate the next token rather than for representation, Muennighoff et al. [24] leverages instruction to fine-tune the model. Therefore, we fine-tune the last hidden states with a low learning rate during the training phase of TRIDENT.
@@
It is important to note that although LLaVA v1.5 may have seen certain images during training, the model is asked to generate textual descriptions of images in an autoregressive manner during training. The textual descriptions focus on the main content of the image, rather than the ”attribute-object” label. Therefore, there is no issue of data leakage when training TRIDENT.
Appendix C Data Statictics
@@
We evaluate our model on three challenging CZSL benchmark datasets: MIT-states [11], C-GQA [25], and VAW-CZSL [38]. MIT-states consists of diverse real-world images labeled by early image search engine technology. C-GQA and VAW-CZSL are two more challenging benchmark datasets that consist of broad collections of in-the-wild images. C-GQA has more one-to-one compositions, while objects in VAW-CZSL share more attributes. Table 6 shows detailed data statistics following the common data splits of MIT-States [11], C-GQA [25] and VAW-CZSL [38]. MIT-States contains 53753 images, with 115 attributes and 245 objects. It comprises 1262 seen compositions and 300/400 (validation/test) unseen compositions. C-GQA is a natural image dataset which contains 39298 images, with 413 attributes and 764 objects. It includes 5,592 seen compositions and 1,040/923 (validation/test) unseen compositions. VAW-CZSL is a larger dataset which contains 440 attributes and 541 objects for 238040 images, and it is split into 11175 seen and 2322/2470 unseen compositions for training and validation/testing, respectively.
@@
Appendix D Implementation details
We use NVIDIA PTX 3090 GPU to train all models under the Pytorch framework [31]. Since $\mathcal{L}_{comp}$ leverages image features during training, we use a Batch Normalization, ReLU and 0.3 dropout for Image embedder. We train TRIDENT by Adam optimizer with weight decay 5e-5, learning rates 1.5e-6 for word embedding as well as 2e-4 for other modules on three datasets. We decay the learning rate by 10 at epoch 30 and 40. The temperature variable of cosine similarity $\delta$ is set to 0.05. For weighting coefficients $\gamma_{ortho},\gamma_{comp},\gamma_{attr},and\gamma_{obj}$, we set them to 0.1, 1, 0.5, and 0.5, respectively.
Appendix E Impact of Hyperparameters
To provide more insight into the effect of visual features and label smoothing, we study on the performance of TRIDENT with respect to different numbers of visual features and different label smoothing factors, respectively. Experiments exploring the impact of hyperparameters are conducted on datasets MIT-States and C-GQA.
Impact of the number of visual features. We illustrate the performance of TRIDENT influenced by different numbers of attribute features in Figure . In Figure 4(a), the performance of our model on MIT-States generally improves with the increasing number of visual features, but subsequently declines. This trend is reasonable, as a greater number of Visual features contains more useful information, thereby enhancing the performance. However, the number of useful features is limited; thus, an excessive number of visual features may introduce redundancy and noise, ultimately hampering the performance of the model.
However, in Figure 4(b), as the number of visual features increases, the performance of the model on C-GQA tends to decline overall. This may be attributed to the model’s strong expressive capability in handling composition reasoning. In the low-noise C-GQA dataset, optimal performance can be achieved using only two features. Increasing the number of features, however, results in heightened model complexity without tangible benefits, potentially impairing generalization to unseen compositions. In contrast, the MIT-States dataset exhibits significant noise; thus, while the increase of visual features may introduce more noise, it also necessitates a greater amount of useful information, which can effectively mitigate the impact of the noise.
Impact of the number of label smoothing factor. The label smoothing factor $\alpha$ modulates the extent to which the model’s confidence in seen compositions is attenuated. Figure 5(a) shows that as $\alpha$ increases, the model’s performance on MIT-States initially improves before subsequently declining. This is because if alpha is too small, label smoothing fails to enhance generalization, while if alpha is too large, it adversely affects the model’s ability to learn the representation of the original labels, resulting in more losses than gains. However, as shown in Figure 5(b), the model achieves the best performance with C-GQA a smaller $\alpha$. This may be attributed to the fact that, compared to everyday objects, LLMs are less familiar with in-the-wild objects, leading to relatively lower quality in the generated auxiliary attributes; thus, a smaller smoothing factor can mitigate the impact.

### v2

@@ 4.1 Experiment Setup
Baselines. We compare our TRIDENT with recent and prominent approaches in CZSL: SymNet Li et al. (2020), CompCos Mancini et al. (2021), Co-CGE Mancini et al. (2022), SCEN Li et al. (2022), OADis Saini et al. (2022), INV Zhang et al. (2022), CANet Wang et al. (2023), and ProCC Huo et al. (2024). We replace their visual backbone with ViT-Large-Patch14-336px and retrain all models with the same number of epochs for the sake of fairness. In addition, although comparing TRIDENT with CLIP-based methods, which rely on the dual-tower architecture, is very unfair due to inadvertent exposure to unseen compositions, we still choose CLIP Radford et al. (2021) and CoOp Zhou et al. (2022) as baselines for their strong zero-shot abilities.
(a) image-to-text retrieval.
(b) text-to-image retrieval (successful cases).
(c) text-to-image retrieval (failure cases).
Figure 3: Qualitative analysis. (a) Top-5 image-to-text retrieval. The first two rows display successful cases, while the last row presents failure cases. For each image, the top row shows the ground-truth, followed by five rows of top-5 predictions. (b) Successful cases and (c) failure cases of top-5 text-to-image retrieval. In all cases, the successful and failure results are tagged in green and red, respectively.
4.2 Results and Discussion
In this section, we compare TRIDENT with state-of-the-art methods. As shown in Table 2, our model surpasses other models by a substantial margin in general. TRIDENT boosts $AUC$ from 13.6%, 5.7%, and 6.7% of the previous state-of-the-art method CANet to the new state-of-the-art performance of 14.2%, 8.0%, and 8.3% with 0.6%, 2.3%, and 1.6% improvement on three datasets, respectively. In addition, TRIDENT achieves 30.9%, 22.6%, 23.4% on the metrics of $HM$, providing 1.1%, 3.7%, and 2.4% improvement on CANet. For MIT-States, our model achieves competitive performance, despite considerable label noise Atzmon et al. (2020). The largest improvement is observed on the $Unseen$ metric, indicating that attribute smoothing helps enhance the generalization ability of the model. We also observe that TRIDENT performs significantly better than CANet regarding all metrics on two more challenging and low-noise datasets C-GQA and VAW-CZSL, indicating the efficacy of our approach. These improvements arise from the utilization of MLLM embeddings and attribute smoothing, which enhance attribute-object disentanglement and consequently facilitate the recognition of unseen compositions while maintaining performance on seen compositions.
4.3 Ablation Study
We ablate the components of TRIDENT on MIT-States to evaluate their contributions. From the ablation results in Table 3, we gain the following observations.
1) TRIDENT outperforms the models without using patch and [CLS] features, indicating that both patch and [CLS] features are crucial, with patch features contributing more.
2) Both w/o FAAs and w/o condition_masks models perform worse than TRIDENT, which validates the importance of filtering out the background noise and extracting the multi-granularity features, respectively.
3) TRIDENT surpasses w/o word_expanding model and w/o attribute_smoothing model on the $Unseen$ metric, yet falls short of them on the $Seen$ metric. The difference between our model and the w/o word_expanding model stems from its more thorough disentanglement, which enhances the recognition of unseen compositions while weakening the identification of seen ones. The disparity between our model and the w/o attribute_smoothing model arises from attribute smoothing, which diminishes the overconfidence of the model in seen compositions, facilitating its generalization to unseen ones. However, the improvement of our model over these two models on $AUC$ and $HM$ indicates the effectiveness of the word expanding and label smoothing strategy.
4) TRIDENT outperforms w/o $\mathcal{L}_{attr}+\mathcal{L}_{obj}$ model on all metrics, confirming that the attribute-object disentanglement module is highly advantageous for generalization from seen compositions to unseen compositions.
5) w/o $\mathcal{L}_{ortho}$ model is inferior to TRIDENT, which suggests that the designed orthogonal regularization is helpful to guarantee that different features extract different visual information.
Impact of word embeddings. Our work leverages the last hidden states of LLaVA v1.5 as word embeddings, while Word2Vec Mikolov (2013) and FastText Bojanowski et al. (2017) are the common word embeddings used for MIT-States in previous works. In Table 4, based on three models: SCEN Li et al. (2022), CANet Wang et al. (2023) and TRIDENT, we compare the performance of using the last hidden states of LLaVA v1.5 (${LLaVA}_{lhs}$) and the sum of Word2Vec and FastText (ft+w2v), respectively. The results indicate that the last hidden states of MLLM capture more complex multimodal semantic information than ordinary word embeddings.
@@ 4.4 Qualitative Analysis
In this section, we use TRIDENT to conduct both image-to-text retrieval and text-to-image retrieval experiments on the three datasets. We first consider image-to-text retrieval, shown in Figure 3(a). For successful cases, such as the image labeled burnt house, we notice that the top four predictions can describe logs burning on a fire in the image. In terms of the image labeled green leaf, another successful case, all predicted attributes can describe the leaf, which is due to attribute smoothing learning more attributes for an object. For the failure cases, such as the image labeled colorful pillow, the prediction of orange pillow can also describe the image.
We then consider text-to-image retrieval. Successful cases are shown in Figure 3(b), while failure cases are shown in Figure 3(c). We observe that all retrieved images of peeled orange are correct. However, the retrieved images of green grapes are all wrong. This is due to the fact that the training images of green grapes are almost entirely filled with a single grape, which makes it difficult for the model to capture the contour features of a bunch of green grapes.
@@ 5 Conclusion
In this work, we propose a novel framework termed TRIDENT to address the challenging CZSL task. First, we leverage feature adaptive aggregation modules to mitigate the impact of background, and utilize learnable condition masks to capture multi-granularity features for attribute-object disentanglement. In addition, we exploit the last hidden states of MLLM to replace ordinary word embeddings, as they can capture complex multimodal semantic information. Moreover, we leverage LLM to generate auxiliary attributes and perform attribute smoothing to diminish overconfidence of the model in seen compositions, which enables it to generalize to unseen compositions more effectively. Extensive experiments conducted on three challenging datasets demonstrate the effectiveness of our method.
Acknowledgments
Our work was supported by the Beijing Natural Science Foundation (No. 4242046).
@@ Appendix A Auxiliary Attributes Generation by LLM and Post-process
As mentioned before, we leverage GPT-3.5 OpenAI [2023] to generate auxiliary attributes for attribute smoothing. The auxiliary attributes are generated based on the composition instead of the object, such as ripe apple and sliced apple, which takes into account the unique contextuality that arises from the combination of attribute and object Zhang et al. [2024]. This ensures that the auxiliary attributes accurately capture the nuances of the specific composition, rather than generating general object-level attributes. Since the generation ability of LLMs is affected by many factors, in this section, we first explore the impact of different prompts. Then we study the influence of the number of auxiliary attributes $t$. In addition, we introduce post-processing of the generated text to eliminate potential noise.
Impact of prompt input to LLM. Users typically interact with LLMs through prompts to request answers to questions, generate text, complete tasks, and more. The model generates text based on the provided prompt, striving to meet the users’ requirements Sahoo et al. [2024]. Therefore, the good design of prompt is significant for eliciting knowledge of LLMs, which enables them to better follow our instructions and generate auxiliary attributes with high quality. We first design prompts with different styles, then input them into GPT-3.5 and observe the quality of generated auxiliary attributes. Some prompt examples for the composition browned beef and ancient building are shown in Table 5 and Table 6, respectively.
@@
As shown in Table 5, the prompt without ”the visual feature of …” may cause the model to produce adjectives that are not specific but generic, such as Savory and Rich. In both Table 5 and Table 6, the prompts starting with ”Now I give you…”, compared to those starting with ”Please give me …”, result in a weaker instruction following ability of the model. Therefore, we choose the prompt: ”Please give me $t$ adjectives that can describe the visual feature of a photo of a/an … well.”
Additionally, we incorporate supplementary instructions to constrain the format of the generated text. The complete prompt is as follows: “Please give me $t$ adjectives that can describe the visual feature of a photo of a/an … well. Please strictly follow the format: give me just $t$ words separated by commas and do not say more.”
@@
Impact of the number of auxiliary attributes $t$. In Table 5, we observe that the generated attributes describe the compositions to varying degrees, with later items in the sequence being less relevant generally. Therefore, we study the influence of the number of auxiliary attributes $t$.
@@
Table 7 and Table 8 show the generated text using different $t$ with compositions large garden and young girl. The results demonstrate that the greater the number, the more generic adjectives with irrelevant information are included, for example, Captivating is generated for both compositions. In addition, with $t$ increasing, the noise in the generated text grows due to the uncertainty of the model about the given composition. For example, the young girl may not be happy, yet the model fails to find ten words to describe her, so it has to guess. Therefore, we set $t$ to 3, which minimizes the general adjectives and noise while retaining useful information.
@@
Post-processing of generated text. GPT-3.5 generates a segment of text, which we need to process into multiple useful words by exploiting regular expressions. However, the auxiliary attributes generated by LLM may contain the attribute of the input composition, for example, generating ancient for ancient building. At this point, we reuse the model to generate $t+1$ adjectives for this composition and select $t$ adjectives that are not the attribute of the input composition.
Appendix B Auxiliary Attribute Quality Assessment
In this section, we evaluate the quality of auxiliary attributes in terms of both correctness and diversity on MIT-States. Two approaches are employed: evaluation based on LLM generation and evaluation using CLIP representations.
Evaluation of the generation quality of LLMs is always performed by querying another LLM multiple times and averaging the results. We assess the quality of auxiliary attributes in terms of correctness and diversity by asking GPT-4o OpenAI [2024] with two crafted prompts, respectively: ”Can the word ’[Auxiliary Attribute]’ describe a/an ’[Composition]’? Please answer ’Yes’ or ’No’.” and ”Does the use of ’[Auxiliary Attribute]’ to describe a [Composition] provide additional information? Please answer ’Yes’ or ’No’.” For each auxiliary attribute of each combination, we query three times and calculate the average accuracy of the ”Yes” responses. As observed in Table 9, the auxiliary attributes are of high quality in terms of both correctness and diversity.
Indicator Test 1 Test 2 Test 3 Average
Correctness 93.2% 92.5% 90.8% 92.6%
Diversity 83.5% 82.9% 83.1% 83.2%
Table 9: Evaluation on the auxiliary attributes on MIT-States.
In addition, we use CLIP and ”a photo of” prompt to compare the similarities between the original composition and [auxiliary attribute][object] ([a-a][o]), [random attribute][object] ([r-a][o]), and [auxiliary attribute] ”object” ([a-a]”o”), respectively. In Table 10, the first similarity greater than the last two respectively indicates auxiliary attributes describe objects correctly, and auxiliary attributes are diverse.
Experiment [a-a][o] [r-a][o] [a-a]”o”
Similarity 0.8389 0.7950 0.8197
Table 10: Similarities of different forms of attribute-objects with the original composition by CLIP.
We also visualize text features of CLIP for $t\ (t=3)$ [a-a][o] and the original [a][o] of randomly selected 40 compositions by t-SNE. In Figure 4, every four features of the composition are in the same cluster (verifying correctness) and the points within the cluster are scattered (verifying diversity).
Figure 4: Visualization of auxiliary attribute-objects ([a-a][o]) and original compositions ([a][o]).
Appendix C Obtainment of the Last Hidden States of MLLM
@@
We input the attribute (object) word into LLaVA-v1.5-7B Liu et al. [2024a], which first tokenizes the word into $o$ tokens. These tokens pass through all Transformer blocks in the MLLM, ultimately generating $o$ embeddings of dimension $d_{m}$ after the last block, named the last hidden states. Subsequently, we apply average pooling to these $o$ embeddings of dimension $d_{m}$ to obtain a $d_{m}$-dimensional embedding that represents the attribute. Since the last hidden states are designed to generate the next token rather than for representation, GritLM Muennighoff et al. [2024] leverages instruction to fine-tune the model. Therefore, we fine-tune the last hidden states with a low learning rate during the training phase of TRIDENT.
@@
It is important to note that although LLaVA v1.5 may have seen certain images during training, the model is asked to generate textual descriptions of images in an autoregressive manner during training. The textual descriptions focus on the main content of the image, rather than the ”attribute-object” label.
Appendix D Dataset Introduction
@@
We evaluate our model on three challenging CZSL benchmark datasets: MIT-States Isola et al. [2015], C-GQA Naeem et al. [2021], and VAW-CZSL Saini et al. [2022]. MIT-States consists of diverse real-world images labeled by early image search engine technology. C-GQA and VAW-CZSL are two more challenging benchmark datasets that consist of broad collections of in-the-wild images. C-GQA has more one-to-one compositions, while objects in VAW-CZSL share more attributes. Table 1 shows detailed data statistics following the common data splits of MIT-States Isola et al. [2015], C-GQA Naeem et al. [2021] and VAW-CZSL Saini et al. [2022]. MIT-States contains 53753 images, with 115 attributes and 245 objects. It comprises 1262 seen compositions and 300/400 (validation/test) unseen compositions. C-GQA is a natural image dataset which contains 39298 images, with 413 attributes and 764 objects. It includes 5,592 seen compositions and 1,040/923 (validation/test) unseen compositions. VAW-CZSL is a larger dataset which contains 440 attributes and 541 objects for 238040 images, and it is split into 11175 seen and 2322/2470 unseen compositions for training and validation/testing, respectively.
Appendix E Impact of Hyperparameters
To provide more insight into the effect of visual features and label smoothing, we study the performance of TRIDENT with respect to different numbers of visual features and different label smoothing factors, respectively. Experiments are conducted on two datasets MIT-States and C-GQA.
Impact of the number of visual features. In Figure 5(a), the performance of our model on MIT-States generally improves with the increasing number of visual features, but subsequently declines. This trend is reasonable, as a greater number of visual features contain more useful information, thereby enhancing the performance. However, the number of useful features is limited; thus, an excessive number of visual features may introduce redundancy and noise, ultimately hampering the performance of the model.
However, in Figure 5(b), as the number of visual features increases, the performance of the model on C-GQA tends to decline overall. This may be attributed to the model’s strong expressive capability in handling composition reasoning. In the low-noise C-GQA dataset, optimal performance can be achieved using only two features. Increasing the number of features, however, results in heightened model complexity without tangible benefits, potentially impairing the performance of the model. In contrast, the MIT-States dataset exhibits significant noise; thus, while the increase in the number of visual features may introduce more noise, it also necessitates a greater amount of useful information, which can effectively mitigate the impact of the noise.
@@ Appendix E Impact of Hyperparameters
Impact of the label smoothing factor. The label smoothing factor $\alpha$ modulates the extent to which the model’s confidence in seen compositions is attenuated. Figure 6(a) shows that as $\alpha$ increases, the model’s performance on MIT-States initially improves before subsequently declining. This is because if $\alpha$ is too small, label smoothing fails to enhance generalization, while if alpha is too large, it adversely affects the model’s ability to learn the representation of the original labels, resulting in more losses than gains. However, as shown in Figure 6(b), the model achieves the best performance on C-GQA with a smaller $\alpha$. This may be attributed to the fact that, compared to everyday objects, LLMs are less familiar with in-the-wild objects, leading to relatively lower quality in the generated auxiliary attributes; thus, a smaller smoothing factor can mitigate the impact.
@@
Appendix F Do Auxiliary Attributes Help or Hurt?
Ablation study in the full paper proves attribute smoothing is indeed helpful. In addition, MIT-States has 115 attributes, and the set of auxiliary attributes for the dataset has 621 elements. According to statistics, 543 auxiliary attributes (the majority) do not appear in the original dataset. Smoothing with these auxiliary attributes does not mislead the model into mistakenly classifying images into other attributes in the dataset. In addition, changing the smoothing factor $\alpha$ and the number of auxiliary attributes $t$ can control the degree of smoothness to achieve the optimal performance.